OpenStudy (anonymous):

Maki uses 20 gallons of a solution with an unknown ethanol concentration and 60 gallons of a 12% ethanol solution to make 80 gallons of a 10% solution. The table shows the amount of each solution. What is the value of x, the concentration of the unknown ethanol solution? a. 0.02 b. 0.04 c. 0.05 d. 0.08

OpenStudy (anonymous):

b

OpenStudy (anonymous):

how though?

OpenStudy (anonymous):

i dont know how to explain it, sorry :(

OpenStudy (anonymous):

are you sure it's b?

OpenStudy (anonymous):

im 85% its b

OpenStudy (kropot72):

Amount of ethanol in 20 gallons of unknown concentration = 20x Amount of ethanol in 60 gallons of 12% solution = 60 * 0.12 Amount of ethanol in 80 gallons of 10% solution = 80 * 0.1 Now we can construct an equation as follows: \[\large 20x+60\times0.12=80\times0.1\] which simplifies to: \[\large 20x+7.2=8\]

OpenStudy (kropot72):

Now just solve the last equation to find the value of x.

OpenStudy (anonymous):

so b?